# A Nonlocal Aw-Rascle-Zhang system with linear pressure term

### v1

@@ 4. Approximation of the nonlocal system
In this section, we come back to system (1.5) and study it using the weak entropy solution $M$ obtained in Section 3.
with initial data
(4.2) $\rho(x,0)=\rho_{0}(x),\qquad u(x,0)=u_{0}(x).$
Let $\mathcal{P}_{c}(\mathbb{R})$ be the space of probability measures with compact support, we use the Wasserstein-1 metric to quantify the distance between elements of $\mathcal{P}_{c}(\mathbb{R}),$ therefore we write the following definition.
@@
Remark 4.2.
[…]
In our setting the $\mathcal{W}_{1}$ convergence is equivalent to the weak-* convergence in the sense of measures. In the following, we consider $\mathcal{P}_{c}(\mathbb{R})$ equipped with the Wasserstein-1 metric. In the next definition we state what we mean for weak solution of (4.1).
@@
  • The initial condition $(\rho^{0},P^{0})$ is attained in the following sense, for any $f\in C_{c}^{\infty}(\mathbb{R})$
    $\lim_{t\to 0+}\int_{\mathbb{R}}f(x)\,d\rho(x,t)=\int_{\mathbb{R}}f(x)\,d\rho^{0}(x),\quad\lim_{t\to 0+}\int_{\mathbb{R}}f(x)\,dP(x,t)=\int_{\mathbb{R}}f(x)\,dP^{0}(x).$
@@
Assume (H) and (1.6). Let $\rho^{0}\in\mathcal{P}_{c}(\mathbb{R})$, $u^{0}\in L^{\infty}(d\rho^{0})$ and $P^{0}=\rho^{0}u^{0}$. We construct a unique couple $(\rho,P)$ using the following procedure.
@@
    Let $M$ be the unique entropy solution of (3.28) with initial datum $M^{0}$ and flux $A$.
@@
Then, the couple $(\rho,P)$ is a weak solution of the ARZ system (4.1) with initial data (4.2) in the sense of Definition 4.3. Moreover, we can define $u(\cdot,t)=\frac{dP(\cdot,t)}{d\rho(\cdot,t)}$ that is the Radon-Nikodym derivative of $P(t)$ with respect to $\rho(t)$.
@@
The functions $M^{0}$ and $A$ defined in the step (i) satisfy the hypotheses of Theorem 3.4, indeed, $M^{0}$ is a nondecreasing function by construction, and $\rho^{0}$ is a probability measure with compact support, then it is possible to find an $R^{0}>0$ such that $\text{supp}\{\rho^{0}\}\subset[-R^{0},R^{0}]$. Regarding $A$, it is possible to see that
[…]
as a consequence $A$ as in (4.3) is Lipschitz continuous. Applying Theorem 3.4 we obtain a unique weak entropy solution of the scalar balance problem (3.28), $M\in BV(\mathbb{R}\times[0,T])$ for a fixed $T>0$. To verify that the couple $(\rho,P)$ constructed as in (4.4) satisfies the Definition 4.3 of weak solution, we apply [5, Lemma 2.1], then the convergence result (3.29) gives $\rho\in C([0,T];\mathcal{P}_{c}(\mathbb{R}))$, while (3.30) implies that $P=-\partial_{t}M\in\mathcal{M}(\mathbb{R})$. Since $M\in BV(\mathbb{R}\times[0,T])$ and $A$ is a Lipschitz function, using Theorem B.1, we know that there exists a bounded Borel measurable function $\psi=\psi(x,t)$ such that
@@
$\mathinner{\!\left\lvert\int_{\mathbb{R}}f^{\prime}(x)A(M(x,t))\,dx-\int_{\mathbb{R}}f^{\prime}(x)A(M^{0}(x))\,dx\right\rvert}\leq|A|_{Lip}\mathinner{\!\left\lVert f^{\prime}\right\rVert}_{L^{\infty}}\mathinner{\!\left\lVert M(\cdot,t)-M^{0}\right\rVert}_{L^{1}}\to 0.$
@@
$\leq\mathinner{\!\left\lVert\phi\ast(M-M^{0})\right\rVert}_{L^{\infty}(\mathbb{R})}\mathinner{\!\left\lvert\int_{\mathbb{R}}f(x)\,d\rho(x,t)-\int_{\mathbb{R}}f(x)\,d\rho^{0}(x)\right\rvert}$
$\leq\mathinner{\!\left\lVert\phi\right\rVert}_{L^{\infty}(\mathbb{R})}\mathinner{\!\left\lVert M-M^{0}\right\rVert}_{L^{1}(\mathbb{R})}\mathinner{\!\left\lvert\int_{\mathbb{R}}f(x)\,d\rho(x,t)-\int_{\mathbb{R}}f(x)\,d\rho^{0}(x)\right\rvert}\to 0,$
@@
Thanks to Theorem 4.4, we can conclude that the solution of the ARZ system (4.1) exists and it is unique, moreover it satisfies the following stability estimate.
@@
[…]
@@
Thanks to Remark 4.2 and the stability estimate (3.31), we know that
$\mathcal{W}_{1}(\rho,\mu)$ $=\mathinner{\!\left\lVert M-\tilde{M}\right\rVert}_{L^{1}(\mathbb{R})}$
[…]
In what follows we show that it is possible to obtain a solution of system (4.1) using atomic initial data constructed using the sticky particle Cucker-Smale dynamics (2.1)-(2.5)-(2.6), we can prove the following results. We refer to [23, Proposition 6.4 and Theorem 6.5] respectively for the proofs.
@@
We refer to Proposition 6.4 and Theorem 6.5 of [23] respectively for the proofs.
@@ Appendix A Convolution of measures and functions
In this section we recall the theory about the convolution product between a function and a measure, in particular we want to see when this kind of operation is well defined. We refer to [2] and to [4] for this section. We start with some definitions.
@@
The following convolution product is defined for almost $x$ with respect to Lebesgue measure and
@@ Appendix B Chain Rule for functions of Bounded Variation
Let $\Omega$ be a bounded open subset of $\mathbb{R}^{n}$, and $f\mathrel{\mathop{\ordinarycolon}}\mathbb{R}\to\mathbb{R}$ be a globally Lipschitz continuous function. Consider $u\in BV(\Omega,\mathbb{R})$ and set $v=f\circ u$. The the function $v$ belongs to $BV(\Omega,\mathbb{R})$ and we have the chain rule

### v2

@@ 4. Approximation of the nonlocal system
In this section, we return to the system (1.5) and analyze it using the framework developed in Section 3 for the cumulative function $M$. Specifically, let $\mathcal{P}_{c}(\mathbb{R})$ denote the space of probability measures with compact support, equipped with the Wasserstein-1 metric, which we define below.
@@
In our setting, the $\mathcal{W}_{1}$ convergence is equivalent to the weak-* convergence in the sense of measures. We remark that, if $M$ and $\tilde{M}$ are the cumulative distribution functions of $\rho$ and $\mu$ respectively, then the following identity holds:
(4.1) $\mathcal{W}_{1}(\rho,\mu)=\|M-\tilde{M}\|_{L^{1}(\mathbb{R})}\,.$
Now, let’s consider
endowed with initial data
(4.3) $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho(\cdot,0)=\rho^{0},\qquad u(\cdot,0)=u^{0}.$}
Here we provide the definition of weak solution of the Cauchy problem (4.2)–(4.3).
@@
  • The initial condition $(\rho^{0},P^{0})$ is attained in the following sense:
    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho(0)=\rho^{0}}\,,\qquad\lim_{t\to 0+}\int_{\mathbb{R}}f(x)\,dP(x,t)=\int_{\mathbb{R}}f(x)\,dP^{0}(x)\quad\forall\,f\in C_{c}^{\infty}(\mathbb{R})\,.$
@@
Assume (H) and (1.6). Let $\rho^{0}\in\mathcal{P}_{c}(\mathbb{R})$, $u^{0}\in L^{\infty}(d\rho^{0})$ and $P^{0}=\rho^{0}u^{0}$. We construct a unique pair $(\rho,P)$ using the following procedure.
@@
    Let $M$ be the unique entropy weak solution of (3.28) with initial datum $M^{0}$ and flux $A$.
@@
Then, the pair $(\rho,P)$ is a weak solution of the ARZ system (4.2) with initial data (4.3) in the sense of Definition 4.2. Moreover, we can define $u(\cdot,t)=\frac{dP(\cdot,t)}{d\rho(\cdot,t)}$ that is the Radon-Nikodym derivative of $P(t)$ with respect to $\rho(t)$.
@@
The functions $M^{0}$ and $A$ defined in the step (i) satisfy the hypotheses of Theorem 3.5. Indeed, $M^{0}$ is a nondecreasing function by construction, and $\rho^{0}$ is a probability measure with compact support, then it is possible to find an $R^{0}>0$ such that $\text{supp}\{\rho^{0}\}\subset[-R^{0},R^{0}]$. Regarding $A$, it is possible to see that
$\|\psi^{0}\|_{L^{\infty}(d\rho^{0})}\leq\|u^{0}\|_{L^{\infty}(d\rho^{0})}+\|\omega\|_{L^{\infty}(\mathbb{R})}<+\infty,$
as a consequence $A$ as in (4.4) is Lipschitz continuous. Applying Theorem 3.5 we obtain a unique entropy weak solution of the scalar balance problem (3.28), $M\in BV(\mathbb{R}\times[0,T])$ for a fixed $T>0$.
To verify that the pair $(\rho,P)$ constructed as in (4.5) satisfies the Definition 4.2 of weak solution, we apply [5, Lemma 2.1], then the convergence result (3.29) gives $\rho\in C([0,T];\mathcal{P}_{c}(\mathbb{R}))$, while (3.30) implies that $P=-\partial_{t}M\in\mathcal{M}(\mathbb{R})$. Since $M\in BV(\mathbb{R}\times[0,T])$ and $A$ is a Lipschitz function, using Theorem B.1, we know that there exists a bounded Borel measurable function $\psi=\psi(x,t)$ such that
@@
$\mathinner{\!\left\lvert\int_{\mathbb{R}}f^{\prime}(x)A(M(x,t))\,dx-\int_{\mathbb{R}}f^{\prime}(x)A(M^{0}(x))\,dx\right\rvert}\leq|A|_{Lip}\|f^{\prime}\|_{L^{\infty}}\|M(\cdot,t)-M^{0}\|_{L^{1}}\to 0.$
@@
$\leq\|\phi\ast(M-M^{0})\|_{L^{\infty}(\mathbb{R})}\mathinner{\!\left\lvert\int_{\mathbb{R}}f(x)\,d\rho(x,t)-\int_{\mathbb{R}}f(x)\,d\rho^{0}(x)\right\rvert}$
$\leq\|\phi\|_{L^{\infty}(\mathbb{R})}\|M-M^{0}\|_{L^{1}(\mathbb{R})}\mathinner{\!\left\lvert\int_{\mathbb{R}}f(x)\,d\rho(x,t)-\int_{\mathbb{R}}f(x)\,d\rho^{0}(x)\right\rvert}\to 0,$
@@
Thanks to Theorem 4.3, we can conclude that the solution of the ARZ system (4.2) exists and it is unique; moreover it satisfies the stability estimate established in the following proposition.
@@
(4.6) $\mathcal{W}_{1}(\rho,\mu)\leq\mathcal{W}_{1}(\rho^{0},\mu^{0})e^{2t\|\phi\|_{L^{\infty}(\mathbb{R})}}+\frac{|A-\tilde{A}|_{Lip}}{2\|\phi\|_{L^{\infty}(\mathbb{R})}}(e^{t\|\phi\|_{L^{\infty}(\mathbb{R})}}-1).$
@@
By means of (4.1), the stability estimate (3.31) rewrites as
$\mathcal{W}_{1}(\rho,\mu)$ $=\|M-\tilde{M}\|_{L^{1}(\mathbb{R})}$
$\leq e^{2t\|\phi\|_{L^{\infty}(\mathbb{R})}}\|M^{0}-\tilde{M}^{0}\|_{L^{1}(\mathbb{R})}+\frac{|A-\tilde{A}|_{Lip}}{2\|\phi\|_{L^{\infty}(\mathbb{R})}}(e^{t\|\phi\|_{L^{\infty}(\mathbb{R})}}-1)$
$=\mathcal{W}_{1}(\rho^{0},\mu^{0})e^{2t\|\phi\|_{L^{\infty}(\mathbb{R})}}+\frac{|A-\tilde{A}|_{Lip}}{2\|\phi\|_{L^{\infty}(\mathbb{R})}}(e^{t\|\phi\|_{L^{\infty}(\mathbb{R})}}-1)$
which is (4.6).
In what follows we show that it is possible to obtain a solution of system (4.2) using atomic initial data constructed using the sticky particle Cucker-Smale dynamics (2.1)-(2.6)-(2.7), the following results hold.
@@
The above results correspond to Proposition 6.4 and Theorem 6.5 in [23]. Since their proofs do no require additional consideration, we omit them.
@@ Appendix A Convolution of measures and functions
In this section we recall the theory about the convolution product between a function and a measure, in particular we want to see when this kind of operation is well defined. We refer to [2, 4] for this section. We start with some definitions.
@@
The following convolution product is defined for almost all $x$ with respect to Lebesgue measure and
@@ Appendix B Chain Rule for functions of Bounded Variation
Let $\Omega$ be a bounded open subset of $\mathbb{R}^{n}$, and $f\mathrel{\mathop{\ordinarycolon}}\mathbb{R}\to\mathbb{R}$ be a globally Lipschitz continuous function. Consider $u\in BV(\Omega,\mathbb{R})$ and set $v=f\circ u$. Then the function $v$ belongs to $BV(\Omega,\mathbb{R})$ and we have the chain rule